I am sorry I do not have my book with me. I am not sure I am solving these two equations correctly could you please help?
Solve each of the following systems by subtitution:
1st problem:
5x - 2y = -5
y - 5x = 3
2nd problem:
4x - 12y = 5
-x + 3y = -1
Thank you for your help.

Answer by anraymer(35)

For the first problem:
5x - 2y = -5
y - 5x = 3


Choose one variable and isolate it. I chose 5x because I can substitute between both equations:
y - 5x = 3
-5x = -y + 3
5x = y - 3
Then substitute into the first equation:
5x - 2y = -5
(y - 3) -2y = -5
Now solve for y:
-y - 3 = -5
-y = -2
y = 2
Once you have y, you can now solve for x by plugging into either equation:
y - 5x = 3
(2) - 5x = 3
-5x = 1
x = -1/5